OpenStudy (anonymous):

Martina combined different amounts of water, vinegar, and rubbing alcohol to make 38 milliliters (ml) of a solution for a science experiment. The amount of water was x ml. The amount of vinegar was 8 ml less than that of water, and the amount of rubbing alcohol was half the amount of vinegar. This is represented in the equation below. x + (x – 8) + 1/2 (x – 8) = 38 What was the difference in the amounts of water and rubbing alcohol Martina combined?

OpenStudy (bahrom7893):

Oh so all we gotta do is find x, then we can find x-8 and finally (1/2)(x-8)

OpenStudy (bahrom7893):

x + x - 8 + x/2 - 4 = 38 2x + x/2 = 38 + 8 + 4 5x/2 = 50 5x = 100 x = 20

OpenStudy (bahrom7893):

That's the amount of water: 20 The amount of rubbing alcohol: (1/2)(20-8) = (1/2)(12) = 6 Difference: 20-6 = 14